# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v3

@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
$\mathbb{P}\left(\bigcup_{\ell=1}^{5\eta^{-}k}{\mathcal{E}}_{\ell}\right)$ $\leq\sum_{\ell=1}^{5\eta^{-}k}\mathbb{P}({\mathcal{E}}_{\ell})$ (48)
$\leq\sum_{\ell=1}^{5\eta^{-}k}\exp\left(-\Omega\big(\ell\sqrt{\log n}\cdot\log n\big)\right)$ (49)
@@ III-B Proof of Corollary 1
We first recall the result in [12, Theorem 1], which is stated as follows.
@@
[12, Theorem 1] Fix $\alpha\in(0,1)$ and $\varepsilon>0$, and suppose that $T\geq(1-\alpha)\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$. Then there exists a group testing algorithm returning an estimate $\hat{S}^{\prime}$ such that $\max\{\lvert S\setminus\hat{S}^{\prime}\rvert,\lvert\hat{S}^{\prime}\setminus S\rvert\leq\alpha k$ with probability 1o(1)1-o(1).
@@
We can use the same idea for the SUBSET problem; since the reasoning is all analogous to that of [12], we omit some details and focus mainly on the differences. We delete a fraction of items arbitrarily close to $\alpha$ and mark them as non-defective. Among those remaining, we apply the approximate recovery algorithm from [18] with $\beta=\eta^{-}$ for some $\eta^{-}=o(1)$ to obtain $\hat{S}^{\prime}$. As noted in [12], a result in [29, Appendix B] shows that this algorithm still succeeds with high probability even when the number of defectives is only known within a factor of $1\pm o(1)$. We then apply the SUBSET algorithm (Algorithm 1) using the reduced ground set and the two-sided estimate $\hat{S}^{\prime}$. While we are unable to directly apply Algorithm 1 and Theorem 2, since these assume that the number of defectives is known, we can easily modify the algorithm and proof to show the result still holds when it is only known to within a factor of $1\pm o(1)$. Specifically, letting $K^{\prime}$ be the (random) number of non-deleted defectives and defining $k^{\prime}\coloneqq\mathbb{E}K^{\prime}=(1-\alpha+\xi)k$, we observe the following:
@@
  • We cannot set the Hamming distance condition in Algorithm 1 to $d_{\mathrm{H}}(\hat{S}^{\prime},\bar{S})\leq 3\eta^{-}K^{\prime}$, but we can safely replace $K^{\prime}$ by the upper bound $\overline{k}^{\prime}$.
  • Similarly, the output set size cannot be set to $(1-\eta^{-})K^{\prime}$, but it suffices to use $|\bar{S}|=(1-\eta^{-})\underline{k}^{\prime}$.
@@
Since $\eta^{-}=o(1)$ and $\underline{k}^{\prime}=k^{\prime}(1-o(1))$, the resulting output is a set of size $|\hat{S}|=(1-\alpha+\xi)k(1-o(1))$, which is at least $(1-\alpha)k$ when $n$ is large enough. The same $1-\alpha+\xi$ factor then appears in the number of tests, which gives the desired result since $\xi$ is arbitrarily small. ∎
@@ III-C Proof of Theorem 3 (Converse for SUPERSET)
We use ideas from [16, 30], showing that in the exact recovery case, for $\theta$ close to 1, there exists a large number of masked items with high probability, and the existence of such masked items leads to a high probability of failure. This “high $\theta$” result can then be transferred to any $\theta\in\big(\frac{\log 2}{1+\log 2},1\big)$ by an argument based on “dummy non-defective items”. (Note that for $\theta\in\big(0,\frac{\log 2}{1+\log 2}\big]$ the rate in Theorem 3 is 1, and this converse is well-known and holds even for two-sided approximate recovery [18, 12]).
We use these ideas to find similar high probability bounds on the number of masked items in the approximate recovery setting for rates above $R^{*}$. In particular, we show that there are $\omega(1)$ masked defective items and $\omega(k)$ masked non-defective items with high probability, regardless of test design. Intuitively, this leads to the SUPERSET problem not being possible for rates above $R^{*}$, since the decoder will not be able to differentiate between masked defective and masked non-defective items, meaning that when attempting to form an estimate $\hat{S}$ of $S$ of size $(1+\eta^{+})k$ such that $\hat{S}\supseteq S$, the best the decoder can do is guess a uniformly random subset of the masked items to add to the set of non-masked defective items. Since there are many masked non-defective items, the probability the decoder selects all the masked defective items in the subset is small.
@@
We use the following result adapted from [16], which we then bootstrap to prove our claims for the SUPERSET problem.
@@
(Adapted from [16, Propositions 3.4 and 3.5]) Let $M_{1}$ denote the number of masked defective items, where there are $k=\Theta(n^{\theta})$ defective items present within a population of size $n$. Let $\xi>0$ be arbitrarily small, and fix $\varepsilon>0$ and $\theta^{\prime}$ sufficiently close to $1$ such that $2(1-\theta^{\prime})<\xi<\varepsilon\theta^{\prime}$. Then if $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$ and $T\leq\big(\frac{1}{\log 2}k\log_{2}k\big)(1-\varepsilon)$, it holds that
@@
Defining the event $\mathscr{B}=\{k\leq\lvert S^{\prime}\rvert\leq k+2\sqrt{k}\log n\}$, it follows due to the multiplicative Chernoff bound (see Appendix -A) that $\mathbb{P}(\mathscr{B})=1-o(1)$. Moreover, conditioned on $\lvert S^{\prime}\rvert\geq k$ (which is implied by $\mathscr{B}$), the resulting distribution of $S$ is indeed the combinatorial prior due to the symmetry of the construction. Thus, an event has $o(1)$ probability under the above distribution if and only if it has $o(1)$ probability under the combinatorial prior. Additionally, letting $M_{1}^{\prime}$ is the number of masked items after the first step of the construction [^5] Specifically, we use this terminology to mean the number of defectives that would be masked if $S^{\prime}$ were the defective set. and $M_{1}$ denote the number after the second step, it holds that $M_{1}^{\prime}\geq M_{1}$. This is because converting defective items to non-defective in step 2 can only decrease the number of masked items, since items that were masked by such “converted” items may no longer be masked.
Using $M^{\prime}_{1}\geq M_{1}$ and Lemma 10, setting $\gamma\coloneqq\varepsilon-\frac{\xi}{\theta^{\prime}}>0$, and denoting the total number of masked items between the first and second steps as $M^{\prime}$, we have
@@
$=\underbrace{\sum_{m^{\prime}=\frac{1}{4}k^{\gamma}}^{\frac{n}{k}k^{\frac{\gamma}{2}}}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M^{\prime}_{1}\geq\frac{1}{4}k^{\gamma}\mid M^{\prime}=m^{\prime}\right)}_{\coloneqq{\mathcal{A}}_{1}}+\underbrace{\sum_{m^{\prime}=\frac{n}{k}k^{\frac{\gamma}{2}}+1}^{n}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M_{1}^{\prime}\geq\frac{1}{4}k^{\gamma}\mid M^{\prime}=m^{\prime}\right)}_{\coloneqq{\mathcal{A}}_{2}}.$ (54)
We proceed by showing that ${\mathcal{A}}_{1}=o(1)$, which in turn implies ${\mathcal{A}}_{2}=1-o(1)$. Under the i.i.d. prior, it is known that the posterior distribution of masked items matches the prior [30, Lemma 3.1]. Thus, since the first step of the generation of $S$ is i.i.d. with parameter $q^{\prime}$, it holds that $(M_{1}^{\prime}\mid M^{\prime}=m^{\prime})\sim\mathrm{Bin}(m^{\prime},q^{\prime})$. Substituting the choice of $q^{\prime}$, the mean is given by $\mu^{\prime}_{1}\coloneqq\mathbb{E}[M_{1}^{\prime}\mid M^{\prime}=m^{\prime}]=m^{\prime}\frac{k+\sqrt{k}\log n}{n}\leq k^{\frac{\gamma}{2}}\big(1+o(1)\big)$ when $m^{\prime}\leq\frac{n}{k}k^{\frac{\gamma}{2}}$. Thus, writing $\frac{1}{4}k^{\gamma}$ in the form $(1+\delta_{4})\mu^{\prime}_{1}$ for some $\delta_{4}>0$ implies that $\delta_{4}=\Omega(k^{\frac{\gamma}{2}})$. We thus have from the multiplicative Chernoff bound (Appendix -A) that
${\mathcal{A}}_{1}$ $\leq\sum_{m^{\prime}=\frac{1}{4}k^{\gamma}}^{\frac{n}{k}k^{\frac{\gamma}{2}}}\mathbb{P}(M^{\prime}=m^{\prime})\exp\big(-\big((1+\delta_{4})\log(1+\delta_{4})-\delta_{4}\big)\mu^{\prime}_{1}\big)$ (55)
$=\sum_{m^{\prime}=\frac{1}{4}k^{\gamma}}^{\frac{n}{k}k^{\frac{\gamma}{2}}}\mathbb{P}(M^{\prime}=m^{\prime})\exp(-\Omega(k^{\frac{\gamma}{2}}\log k))$ (56)
$\leq\exp(-\Omega(k^{\frac{\gamma}{2}}\log k))$ (57)
@@
where (56) follows since $1+\delta_{4}\geq\delta_{4}=\Omega(k^{\frac{\gamma}{2}})$, and (57) follows by bounding the sum of the probabilities by 1. This implies that ${\mathcal{A}}_{2}=1-o(1)$, which in turn implies via (54) that
@@
Hence, $M^{\prime}=\omega(\frac{n}{k})$ with high probability. We proceed to show that the same is true for $M$, the number of masked items after the second step. The idea is to exploit the fact that $S^{\prime}$ and $S$ are sufficiently “close together” under the high probability event $\mathscr{B}$.
@@
  • (65) follows by bounding the probability that all $j\in{\mathcal{M}}_{t}^{\prime}(i)$ are converted to non-defective by the probability of this happening for a single $j$ (e.g., the one with the lowest index). We can then bound the probability of this single item being converted by $\frac{2\sqrt{k}\log n}{k}$, since on $\mathscr{B}$ at most $2\sqrt{k}\log n$ such items get chosen uniformly at random from a set of size at least $k$ to be made non-defective;
@@
Dropping the bounded tests-per-item assumption. We now argue that the same finding is true for any test design. Consider an arbitrary test design $\mathsf{X}$}, and partition the items $[n]$ into $({\mathcal{N}}_{\mathrm{low}},{\mathcal{N}}_{\mathrm{high}})$, where ${\mathcal{N}}_{\mathrm{low}}$ is the set of items that appear in at most $\log^{3}n$ tests, and ${\mathcal{N}}_{\mathrm{high}}$ contains the remaining items. We may assume without loss of generality that no test in $\mathsf{X}$} contains more than $n^{1-\theta}\log n$ items, since [16, Lemma 3.7] shows that with probability 1o(1)1-o(1), all tests containing more items are positive (and thus their outcomes could have been guessed reliably without performing them). Under this assumption, a simple degree-counting argument in [16, Lemma 3.8] shows that $\lvert{\mathcal{N}}_{\mathrm{low}}\rvert=n-o(n)$. Additionally, while Lemma 10 is stated for the number of masked defectives, the proof in [16] actually shows this result for masked defectives in ${\mathcal{N}}_{\mathrm{low}}$ (again see [16, Lemma 3.8] and its subsequent steps). In other words, in Lemma 10 we may replace $M_{1}$ by $M_{1}^{\rm low}$, defined as the number of masked defectives in ${\mathcal{N}}_{\mathrm{low}}$. We may then apply steps (52)-(66) with all such quantities (i.e., $M_{1}$, $^{\prime}_{1}, MM^{\prime}$, $M$, and the associated sets $\mathcal{M},\mathcal{M}^{\prime}$) replaced by their counterparts that only count items in ${\mathcal{N}}_{\mathrm{low}}$; apart from this modification, the analysis remains identical. It follows that there are at least $\frac{n}{2k}k^{\frac{\gamma}{2}}$ masked non-defectives in ${\mathcal{N}}_{\mathrm{low}}$ with probability 1o(1)1-o(1), which trivially implies the same lower bound on the total number of masked non-defectives.
@@
The steps in this case are similar to the Case 1, so we only provide an outline while highlighting the key differences. Analogous to Case 1, we initially assume that every item is included in at most $n^{\xi}$ tests for some arbitrarily small $\xi>0$; [^6] This is only different from $\log^{3}n$ because we use a result from [30] instead of [16], and they happened to choose slightly different values. The analysis in [30] could easily be modified to use $\log^{3}n$ instead, but we find it more convenient to apply their results exactly as stated. this assumption is dropped in the same way as above, so we do not repeat the details.
@@
(Adapted from [30, Eqn. 3.12]) For any $\varepsilon\in(0,1)$, when $k=\Theta(n^{\theta})$ for some $\theta\in\big[\frac{1}{2},1\big)$ and $T\leq\big(\frac{1}{\log 2}k\log_{2}\frac{n}{k}\big)(1-\varepsilon)$, the number of masked defectives $M_{1}$ satisfies
@@
$=\sum_{m^{\prime}=\frac{k}{4n}k^{1+\xi}}^{n}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M^{\prime}_{1}\geq\frac{k}{4n}k^{1+\xi}\mid M^{\prime}=m^{\prime}\right)$ (69)
$=\underbrace{\sum_{m^{\prime}=\frac{k}{4n}k^{1+\xi}}^{k^{1+\frac{\xi}{2}}}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M_{1}^{\prime}\geq\frac{k}{4n}k^{1+\xi}\mid M^{\prime}=m^{\prime}\right)}_{\coloneqq{\mathcal{A}}_{3}}$
$+\underbrace{\sum_{m^{\prime}=k^{1+\frac{\xi}{2}}+1}^{n}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M_{1}^{\prime}\geq\frac{k}{4n}k^{1+\xi}\mid M^{\prime}=m^{\prime}\right)}_{\coloneqq{\mathcal{A}}_{4}}.$ (70)
@@
  • Finally, we drop the assumption that each item appears in at most $n^{\xi}$ tests; this is done by noting that Lemma 11 still holds when “number of masked items” is replaced by “number of masked items appearing in at most $n^{\xi}$ tests” (see [30, Procedure 3.1, Step 1]) and using the same reasoning as Case 1.
@@
We have established that there are $\omega(k)$ masked non-defectives with high probability, and at least one masked defective (see Lemmas 10 and 11). We proceed to use this to bound $\mathbb{P}(\mathrm{suc})\coloneqq\mathbb{P}(\hat{S}\supseteq S)$, where $\hat{S}$ is the output of an arbitrary decoding algorithm. To do so, we first recall a fact from [16].

### v4

@@ III-A Proof of Theorem 2 (Achievability for SUBSET)
$\mathbb{P}\left(\bigcup_{\ell=1}^{5\eta k}{\mathcal{E}}_{\ell}\right)$ $\leq\sum_{\ell=1}^{5\eta k}\mathbb{P}({\mathcal{E}}_{\ell})$ (66)
$\leq\sum_{\ell=1}^{5\eta k}\exp\left(-\Omega\big(\ell\sqrt{\log n}\cdot\log n\big)\right)$ (67)
@@ III-B Proof of Corollary 1
We first recall the result in [32, Theorem 1], which is stated as follows.
@@
[32, Theorem 1] Fix $\alpha\in(0,1)$ and $\varepsilon>0$, and suppose that $T\geq(1-\alpha)\big(k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$. Then there exists a group testing algorithm returning an estimate $\hat{S}^{\prime}$ such that $\max\{\lvert S\setminus\hat{S}^{\prime}\rvert,\lvert\hat{S}^{\prime}\setminus S\rvert\}\leq\alpha k$ with probability 1o(1)1-o(1).
@@
We can use the same idea for the SUBSET problem; since the reasoning is all analogous to that of [32], we omit some details and focus mainly on the differences. We delete a fraction of items arbitrarily close to $\alpha$ and mark them as non-defective. Among those remaining, we apply the approximate recovery algorithm from [28] with $\beta=\eta$ for some $\eta=o(1)$ to obtain $\hat{S}^{\prime}$. As noted in [32], a result in [30, Appendix B] shows that this algorithm still succeeds with high probability even when the number of defectives is only known within a factor of $1\pm o(1)$. We then apply the SUBSET algorithm (Algorithm 1) using the reduced ground set and the two-sided estimate $\hat{S}^{\prime}$. While we are unable to directly apply Algorithm 1 and Theorem 2, since these assume that the number of defectives is known, we can easily modify the algorithm and proof to show the result still holds when it is only known to within a factor of $1\pm o(1)$. Specifically, letting $K^{\prime}$ be the (random) number of non-deleted defectives and defining $k^{\prime}\coloneqq\mathbb{E}K^{\prime}=(1-\alpha+\xi)k$, we observe the following:
@@
  • We cannot set the Hamming distance condition in Algorithm 1 to $d_{\mathrm{H}}(\hat{S}^{\prime},\bar{S})\leq 3\eta k^{\prime}$, but we can safely replace $K^{\prime}$ by the upper bound $\overline{k}^{\prime}$.
  • Similarly, the output set size cannot be set to $(1-\eta)K^{\prime}$, but it suffices to use $|\bar{S}|=(1-\eta)\underline{k}^{\prime}$.
@@
Since $\eta=o(1)$ and $\underline{k}^{\prime}=k^{\prime}(1-o(1))$, the resulting output is a set of size $|\hat{S}|=(1-\alpha+\xi)k(1-o(1))$, which is at least $(1-\alpha)k$ when $n$ is large enough. The same $1-\alpha+\xi$ factor then appears in the number of tests, which gives the desired result since $\xi$ is arbitrarily small. ∎
@@ III-C Proof of Theorem 3 (Converse for SUPERSET)
Intuitively, an item being masked means that it does not affect the test outcomes of any tests it is placed in, meaning the tests provide no information about its defectivity status. Thus, the best any decoder can do is guess the more likely of the two possibilities. We use ideas from [10, 5], showing that in the exact recovery case, for $\theta$ close to 1, there exists a large number of masked items with high probability, and the existence of such masked items leads to a high probability of failure. They then transfer this “high $\theta$” result to any $\theta\in\big(\frac{\log 2}{1+\log 2},1\big)$ by an argument based on “dummy non-defective items”, which we will also use. (Note that for $\theta\in\big(0,\frac{\log 2}{1+\log 2}\big]$ the rate in Theorem 3 is 1, and this converse is well-known and holds even for two-sided approximate recovery [28, 32]).
We use these ideas to find similar high probability bounds on the number of masked items, suitably modified to suit our approximate recovery setting, for rates above $R^{*}$ (defined in (7)). In particular, we show that there are $\omega(1)$ masked defective items and $\omega(k)$ masked non-defective items with high probability, regardless of test design. Intuitively, this will lead to a converse for SUPERSET because when attempting to form an estimate $\hat{S}$ of $S$ of size $(1+\eta)k$ such that $\hat{S}\supseteq S$, the decoder will be unable to distinguish masked defectives from masked non-defectives. Since there are many more of the latter, any decoder will have a low probability of choosing all masked defectives in its estimate.
@@
We use the following result adapted from [10], which we then use to prove our claims for the SUPERSET problem.
@@
(Adapted from [10, Propositions 3.4 and 3.5]) Let $M_{1}$ denote the number of masked defective items, where there are $k=\Theta(n^{\theta})$ defective items present within a population of size $n$. Let $\xi>0$ be arbitrarily small, and fix $\varepsilon>0$ and $\theta^{\prime}$ sufficiently close to $1$ such that $2(1-\theta^{\prime})<\xi<\varepsilon\theta^{\prime}$. Then if $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$ and $T\leq\big(\frac{1}{\log 2}k\log_{2}k\big)(1-\varepsilon)$, it holds that
@@
Defining the event $\mathscr{B}=\{k\leq\lvert S^{\prime}\rvert\leq k+2\sqrt{k}\log n\}$, it follows from the multiplicative Chernoff bound (see Appendix -A) that $\mathbb{P}(\mathscr{B})=1-o(1)$. Moreover, conditioned on $\lvert S^{\prime}\rvert\geq k$ (which is implied by $\mathscr{B}$), the resulting distribution of $S$ is indeed the combinatorial prior due to the symmetry of the construction. Thus, an event has $o(1)$ probability under the above distribution if and only if it has $o(1)$ probability under the combinatorial prior. Additionally, letting $M_{1}^{\prime}$ be the number of masked defectives after the first step of the construction (i.e., the number of defectives that would be masked if $S^{\prime}$ were the defective set) and $M_{1}$ denote the number after the second step, it holds that $M_{1}^{\prime}\geq M_{1}$. This is because converting defective items to non-defective in step 2 can only decrease the number of masked items (and, in turn, masked defectives), since items that were masked by such “converted” items may no longer be masked.
Using $M^{\prime}_{1}\geq M_{1}$ and Lemma 10, setting $\gamma\coloneqq\varepsilon-\frac{\xi}{\theta^{\prime}}>0$, and denoting the total number of masked items after the first step of the construction by $M^{\prime}$ (i.e., the analog of $M_{1}^{\prime}$ but with both defectives and non-defectives), we have
@@
$=\underbrace{\sum_{m^{\prime}=\frac{1}{4}k^{\gamma}}^{\frac{n}{k}k^{\frac{\gamma}{2}}}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M^{\prime}_{1}\geq\frac{1}{4}k^{\gamma}\mid M^{\prime}=m^{\prime}\right)}_{\coloneqq{\mathcal{A}}_{1}}$
$+\underbrace{\sum_{m^{\prime}=\frac{n}{k}k^{\frac{\gamma}{2}}+1}^{n}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M_{1}^{\prime}\geq\frac{1}{4}k^{\gamma}\mid M^{\prime}=m^{\prime}\right)}_{\coloneqq{\mathcal{A}}_{2}}.$ (72)
We proceed by showing that ${\mathcal{A}}_{1}=o(1)$, which in turn implies ${\mathcal{A}}_{2}=1-o(1)$. Under the i.i.d. prior, it is known that the posterior probability of being defective for each masked item matches its prior [5, Lemma 3.1], which can be shown via Bayes’ rule and the fact that a single masked item’s defectivity status has no impact on any test outcomes. Since the first step of the generation of $S$ is i.i.d. with parameter $q^{\prime}$, it holds that any masked item following this first step has posterior probability $q^{\prime}$ of being defective. As a result, this implies that $\mu^{\prime}_{1}(m^{\prime})\coloneqq\mathbb{E}[M_{1}^{\prime}\mid M^{\prime}=m^{\prime}]=m^{\prime}\frac{k+\sqrt{k}\log n}{n}\leq k^{\frac{\gamma}{2}}\big(1+o(1)\big)$ when $m^{\prime}\leq\frac{n}{k}k^{\frac{\gamma}{2}}$. Thus, Markov’s inequality implies that
${\mathcal{A}}_{1}$ $\leq\sum_{m^{\prime}=\frac{1}{4}k^{\gamma}}^{\frac{n}{k}k^{\frac{\gamma}{2}}}\mathbb{P}(M^{\prime}=m^{\prime})\frac{\mu_{1}^{\prime}(m^{\prime})}{\frac{1}{4}k^{\gamma}}$ (73)
$\leq\sum_{m^{\prime}=\frac{1}{4}k^{\gamma}}^{\frac{n}{k}k^{\frac{\gamma}{2}}}\mathbb{P}(M^{\prime}=m^{\prime})\frac{4k^{\frac{\gamma}{2}}(1+o(1))}{k^{\gamma}}$ (74)
$\leq 4k^{-\frac{\gamma}{2}}(1+o(1))$ (75)
@@
where (74) follows from the above bound on $\mu_{1}^{\prime}(m^{\prime})$ for $m^{\prime}\leq\frac{n}{k}k^{\frac{\gamma}{2}}$, and (75) follows by bounding the sum of the probabilities by 1. This implies that ${\mathcal{A}}_{2}=1-o(1)$, which in turn implies via (72) that
@@
Hence, $M^{\prime}=\omega(\frac{n}{k})$ with probability approaching one. We proceed to show that the same is true for $M$, the number of masked items after the second step. The idea is to exploit the fact that $S^{\prime}$ and $S$ are sufficiently “close together” under the high probability event $\mathscr{B}$.
@@
  • (83) follows by bounding the probability that all $j\in{\mathcal{M}}_{t}^{\prime}(i)$ are converted to non-defective by the probability of this happening for a single $j$ (e.g., the one with the lowest index). We can then bound the probability of this single item being converted by $\frac{2\sqrt{k}\log n}{k}$, since under the event $\mathscr{B}$ at most $2\sqrt{k}\log n$ such items get chosen uniformly at random from a set of size at least $k$ to be made non-defective;
@@
Dropping the bounded tests-per-item assumption. We now argue that the same finding is true for any test design. To do so, we require the following results adapted from [10, 5].
Lemma 11.
(Adapted from [5, Lemma 3.3] and [10, Lemma 3.8]) Let $\Gamma=\frac{n}{k}\log n$, and let ${\mathcal{I}}_{t}\subseteq[n]$ be the set of items included in test $t$. Then, if $T=O(k\log n)$:
  • (i)
    It holds for any test design $\mathsf{X}$} that
    $\hskip-3.0pt\mathbb{P}(\exists t\in\{1,2,\dots,T\}:\lvert{\mathcal{I}}_{t}\rvert>\Gamma\wedge\lvert{\mathcal{I}}_{t}\cap S\rvert=0)=o(1).$ (85)
    That is, it holds with probability 1o(1)1-o(1) that every test containing more than $\Gamma$ items is positive.
  • (ii)
    If all tests in $\mathsf{X}$} contain at most $\Gamma$ items, then there are no(n)n-o(n) items that appear in at most $\log^{3}n$ tests.
We provide the proof of this lemma in Appendix -C. Now, consider an arbitrary test design $\mathsf{X}$}, and partition the items $[n]$ into $({\mathcal{N}}_{\mathrm{low}},{\mathcal{N}}_{\mathrm{high}})$, where ${\mathcal{N}}_{\mathrm{low}}$ is the set of items that appear in at most $\log^{3}n$ tests, and ${\mathcal{N}}_{\mathrm{high}}$ contains the remaining items. By the first part of Lemma 11, we may assume without loss of generality that no test in $\mathsf{X}$} contains more than $\frac{n}{k}\log n$ items, since with probability 1o(1)1-o(1), all tests containing more items are positive (and thus their outcomes could have been guessed reliably without performing them). Under this assumption, the second part of Lemma 11 implies that $\lvert{\mathcal{N}}_{\mathrm{low}}\rvert=n-o(n)$.
The key observation is that while Lemma 10 is stated for the number of masked defectives, the proof in [10] (under the condition $\lvert{\mathcal{N}}_{\mathrm{low}}\rvert=n-o(n)$) actually shows this result for masked defectives in ${\mathcal{N}}_{\mathrm{low}}$; this can directly be seen from the first step after Lemma 3.8 in [10]. In other words, Lemma 10 still remains true when we replace $M_{1}$ by $M_{1}^{\rm low}$, defined as the number of masked defectives in ${\mathcal{N}}_{\mathrm{low}}$. We may then apply steps (70)-(84) with all such quantities (i.e., $M_{1}$, $^{\prime}_{1}, MM^{\prime}$, $M$, and the associated sets $\mathcal{M},\mathcal{M}^{\prime}$) replaced by their counterparts that only count items in ${\mathcal{N}}_{\mathrm{low}}$; apart from this modification, the analysis remains identical. It follows that there are at least $\frac{n}{2k}k^{\frac{\gamma}{2}}$ masked non-defectives in ${\mathcal{N}}_{\mathrm{low}}$ with probability 1o(1)1-o(1), which trivially implies the same lower bound on the total number of masked non-defectives.
@@
The steps in this case are similar to Case 1, so we only provide an outline while highlighting the key differences. Analogous to Case 1, we initially assume that every item is included in at most $n^{\xi}$ tests for some arbitrarily small $\xi>0$; [^6] This is only different from $\log^{3}n$ because we use a result from [5] instead of [10], and they happened to choose slightly different values. The analysis in [5] could easily be modified to use $\log^{3}n$ instead, but we find it more convenient to apply their results exactly as stated. this assumption will later be dropped in the same way as above.
@@
(Adapted from [5, Eqn. 3.12]) For any $\varepsilon\in(0,1)$, when $k=\Theta(n^{\theta})$ for some $\theta\in\big[\frac{1}{2},1\big)$ and $T\leq\big(\frac{1}{\log 2}k\log_{2}\frac{n}{k}\big)(1-\varepsilon)$, the number of masked defectives $M_{1}$ satisfies
@@
$\hskip-5.9pt=\sum_{m^{\prime}=\frac{k}{4n}k^{1+\xi}}^{n}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M^{\prime}_{1}\geq\frac{k}{4n}k^{1+\xi}\mid M^{\prime}=m^{\prime}\right)$ (88)
$\hskip-5.9pt=\underbrace{\sum_{m^{\prime}=\frac{k}{4n}k^{1+\xi}}^{k^{1+\frac{\xi}{2}}}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M_{1}^{\prime}\geq\frac{k}{4n}k^{1+\xi}\mid M^{\prime}=m^{\prime}\right)}_{\coloneqq{\mathcal{A}}_{3}}$
$\hskip-5.9pt+\underbrace{\sum_{m^{\prime}=k^{1+\frac{\xi}{2}}+1}^{n}\mathbb{P}(M^{\prime}=m^{\prime})\mathbb{P}\left(M_{1}^{\prime}\geq\frac{k}{4n}k^{1+\xi}\mid M^{\prime}=m^{\prime}\right)}_{\coloneqq{\mathcal{A}}_{4}}$ (89)
@@
  • Finally, we drop the assumption that each item appears in at most $n^{\xi}$ tests; this is done by noting that Lemma 12 still holds when “number of masked items” is replaced by “number of masked items appearing in at most $n^{\xi}$ tests” (see [5, Procedure 3.1, Step 1]) and using the same reasoning as Case 1.
@@
We have established that there are $\omega(k)$ masked non-defectives with high probability, and at least one masked defective (see Lemmas 10 and 12). We proceed to use this to bound $\mathbb{P}(\mathrm{suc})\coloneqq\mathbb{P}(\hat{S}\supseteq S)$, where $\hat{S}$ is the output of an arbitrary decoding algorithm. To do so, we first recall a fact from [10].